# The $\infty$-S test via quantile affine LASSO

### v2

@@ 5.5 Robustness of the level to non-Gaussian errors
Under the null $H_{0}:A{\boldsymbol{\beta}}={\boldsymbol{b}}$ and for a nominal level $\alpha=0.01$, we show the effective level of the $F$-test and the $\infty$-S test as a function of the degrees of freedom for Student errors d.f. $\in\{1,2,3,4,5,10,100\}$. We choose $n=100$, $p=20$, $A$ is the first $m=5$ rows of the finite difference matrix of Section 5.4 and ${\boldsymbol{\beta}}$ are $p$ samples from the standard Gaussian distribution; $\boldsymbol{b}$ is calculated as $A{\boldsymbol{\beta}}$. In a robustness analysis of the F-test, Ali and Sharma (1996) observe that major determinant of the sensitivity to nonnormality of the errors is the extent of the nonnormality of the regressors or the extent of presence of ‘leveraged’ (influential) observations. So we simulate the entries of the matrix $X$ as a sample from $T$, where $T$ follows a student distribution with 2 degrees of freedom.
@@ 6 Conclusions
The $\infty$-Sτ test allows to test general linear null hypotheses for linear models and for any quantile $\tau$. The test function is simple to implement and can handle high-dimensional problems. We emphasize the $\infty$-S test which corresponds to the $\ell_{1}$-LASSO penalty, and showed that other norms than the $\ell_{1}$-norm could be employed, leading to other tests. Sardy et al. (2022) show that using $\ell_{1}$ leads to high power under sparse alternatives (that is, most entries, but not all, of $A\boldsymbol{\beta}-\boldsymbol{b}$ are zero when $H_{1}$ is true), and, under dense alternatives (that is, most or all entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ are non-zero), that the sign test based on $\ell_{2}$ has more power than the $\infty$-S test. To get nearly the best power with a single test regardless whether the alternative hypothesis is sparse or dense, Sardy et al. (2022) propose the $\oplus$-test.
The $\infty$-S test can be used to derive confidence regions that are dual to tests. The $(1-\alpha)$-confidence region for $A\boldsymbol{\beta}$ based on the $\infty$-S test consists of all vectors $\boldsymbol{b}$ that are not rejected for $H_{0}:A\boldsymbol{\beta}=\boldsymbol{b}$. In particular, to test $H_{0}:\ \beta_{j}=b_{j}$ for some fixed $j\in\{1,\ldots,p\}$, one uses $A=\boldsymbol{e}_{j}^{\rm T}$, the $j$th canonical vector, and the calculated statistic is $s=|\boldsymbol{x}_{j}^{\rm T}\boldsymbol{\omega}|$, where $\boldsymbol{\omega}={\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j})$. So the $(1-\alpha)$-confidence interval for $\beta_{j}$ is the set of all $b_{j}$ such that $|\boldsymbol{x}_{j}^{\rm T}{\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j}(b_{j}))|\leq c_{\alpha}$, where $\hat{\boldsymbol{\beta}}_{-j}(b_{j})\in\arg\min_{\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p-1}}\|\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\tilde{{\boldsymbol{\beta}}}\|_{1}$ and $X_{-j}$ is the $X$ data matrix without its $j$th column.
@@
7 Acknowledgement
This work was partially supported by the Swiss National Science Foundation grant 200021E_213166 and the Academic Rising Star Project No. 2024-YJRC-LXY-005 of the National University of Defense and Technology, China.
@@ Appendix A Proof of Theorem 1
Consider the primal cost function in (5). This primal cost being convex, the gap between the primal and dual cost functions is zero at optimality. The dual cost can be derived as follows
@@
where $\gamma()$ applied componentwise to the penalized vector ${\boldsymbol{\kappa}}=A\boldsymbol{\beta}^{\star}-\boldsymbol{b}$ is defined by $\gamma(\kappa_{k})={\rm sign}(\kappa_{k})$ if $\kappa_{k}\neq 0$, and $\gamma(0)\in[-1,1]$, for $k=1,\ldots,m$. We are interested in finding the smallest $\lambda$ for the solution to (5) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and (15) is
@@
But $X^{\rm T}{\boldsymbol{\omega}}\perp{\rm ker}(A)$ and $A$ is full row rank, so there exists a unique ${\boldsymbol{\alpha}}\in{\mathbb{R}}^{m}$ such that $X^{\rm T}{\boldsymbol{\omega}}=A^{\rm T}{\boldsymbol{\alpha}}$; moreover ${\boldsymbol{\alpha}}=(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}$ is this unique solution. The smallest $\lambda$ satisfying conditions (16) is $\|{\boldsymbol{\alpha}}\|_{\infty}$, so
$\lambda_{0}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty}.$
Under $H_{0}$, the dual cost is $g({\boldsymbol{\omega}}^{\star}):=-{{\boldsymbol{\omega}}^{\star}}^{\rm T}\boldsymbol{y}+0+{{\boldsymbol{\omega}}^{\star}}^{\rm T}X\boldsymbol{\beta}^{\star}$, and the primal cost is $f(\boldsymbol{\beta}^{\star}):=\|\boldsymbol{y}-X\boldsymbol{\beta}^{\star}\|_{1}$ with $\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}_{H_{0}}$. Since the duality gap is zero, then $g({\boldsymbol{\omega}}^{\star})=f(\boldsymbol{\beta}^{\star})$, that is, $-{{\boldsymbol{\omega}}^{\star}}^{\rm T}(\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}})=\|\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}\|_{1}$. Consequently ${\boldsymbol{\omega}}^{\star}=\gamma(X\hat{\boldsymbol{\beta}}_{H_{0}}-{\boldsymbol{y}})$, with the same definition of $\gamma()$ as above. Some entries {${{\cal I}\in\{1,\ldots,n\}}$ of the residuals ${\boldsymbol{r}}={\boldsymbol{y}}-X\hat{\boldsymbol{\beta}}_{H_{0}}$ are different from zero, in which case ${\boldsymbol{\omega}}^{\star}_{\cal I}={\rm sign}(-{\boldsymbol{r}}_{I})$. Let now $K\in{\mathbb{R}}^{p\times(p-m)}$ be a basis for the kernel of $A$. We have that $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}={\boldsymbol{0}}_{p-m}$. Letting $X_{\cal I}$ be the rows of $X$ which indexes are in ${\cal I}$, $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}^{\star}=K^{\rm T}(X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}+X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}})\equiv 0$ iff $K^{\rm T}X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}=-K^{\rm T}X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}$. Solving this linear system leads to the remaining vector of the dual ${\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}\in{\mathbb{R}}^{p-m}$.
@@ Appendix C Proof of Theorem 4
It is well known that the LAD optimization problem (here with $\boldsymbol{y}=\boldsymbol{y}_{-A}$ and $X=X_{-A}$) can be rewritten as the linear program:
$\min_{\boldsymbol{r}_{+}\geq 0,\boldsymbol{r}_{-}\geq 0,\boldsymbol{\beta}}(\boldsymbol{r}_{+}^{\rm T},\boldsymbol{r}_{-}^{\rm T})^{\rm T}\boldsymbol{1}_{2n}\quad{\rm s.t.}\quad\boldsymbol{y}_{-A}=X_{-A}\boldsymbol{\beta}_{-A}+\boldsymbol{r}_{+}-\boldsymbol{r}_{-},$
where $\boldsymbol{1}_{2n}$ is the vector of ones of length $2n$. Using the Lagrange multiplier dual variable $\boldsymbol{\omega}$, the dual can be derived as follows.
@@
with $w_{i}=1=$ when $(r_{+})_{i}>0$, $w_{i}=-1$ when $(r_{-})_{i}>0$, so $\boldsymbol{\omega}$ is the sign of the residuals $\boldsymbol{y}_{-A}-X_{-A}\hat{{\boldsymbol{\beta}}}_{-A}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}$ when the residuals are non-zero. So the dual problems of the constrained LAD and affine LAD LASSO when $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ are the same.
@@ Appendix D Proof of Theorem 5
The proof follows the same lines as the proof of Theorem 1. The dual cost to (12) is the same except that the constraint on ${\boldsymbol{\omega}}\in[-\tau,1-\tau]^{n}$ with the same KKT conditions. We are interested in finding the smallest $\lambda$ for the solution to (12) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}^{\tau}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and the KKT conditions are
@@
But $X^{\rm T}{\boldsymbol{\omega}}\perp{\rm ker}(A)$ and $A$ is full row rank, so there exists a unique ${\boldsymbol{\alpha}}\in{\mathbb{R}}^{m}$ such that $X^{\rm T}{\boldsymbol{\omega}}=A^{\rm T}{\boldsymbol{\alpha}}$; moreover ${\boldsymbol{\alpha}}=(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}$ is this unique solution. The smallest $\lambda$ satisfying conditions (17) is $\|{\boldsymbol{\alpha}}\|_{\infty}$, so
$\lambda_{0}^{\tau}(\boldsymbol{y},X,A)=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}\|_{\infty}.$
Under $H_{0}$, the dual cost is $g({\boldsymbol{\omega}}^{\star}):=-{{\boldsymbol{\omega}}^{\star}}^{\rm T}\boldsymbol{y}+0+{{\boldsymbol{\omega}}^{\star}}^{\rm T}X\boldsymbol{\beta}^{\star}$, and the primal cost is $f(\boldsymbol{\beta}^{\star}):=\|\boldsymbol{y}-X\boldsymbol{\beta}^{\star}\|_{\rho_{\tau}}$ with $\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}$. Since the duality gap is zero, then $g({\boldsymbol{\omega}}^{\star})=f(\boldsymbol{\beta}^{\star})$, that is, $-{{\boldsymbol{\omega}}^{\star}}^{\rm T}{\boldsymbol{r}}=\|\boldsymbol{r}\|_{\rho_{\tau}}$ with the residuals ${\boldsymbol{r}}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}$. Consequently ${\boldsymbol{\omega}}^{\star}=\gamma_{\tau}({\bf r})$ with $\gamma_{\tau}(r_{k})=1-\tau$ if $r_{k}<0$, $\gamma_{\tau}(r_{k})=-\tau$ if $r_{k}>0$ and $\gamma_{\tau}(0)\in[\tau-1,\tau]$, for $k=1,\ldots,m$. Some entries {${{\cal I}\in\{1,\ldots,n\}}$ of the residuals ${\boldsymbol{r}}$ are different from zero, in which case ${\boldsymbol{\omega}}^{\star}_{\cal I}=\gamma_{\tau}({\boldsymbol{r}}_{I})$. Let now $K\in{\mathbb{R}}^{p\times(p-m)}$ be a basis for the kernel of $A$. We have that $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}={\boldsymbol{0}}_{p-m}$. Letting $X_{\cal I}$ be the rows of $X$ which indexes are in ${\cal I}$, $K^{\rm T}X^{\rm T}{\boldsymbol{\omega}}^{\star}=K^{\rm T}(X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}+X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}})\equiv 0$ iff $K^{\rm T}X_{{\cal I}^{c}}^{\rm T}{\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}=-K^{\rm T}X_{\cal I}^{\rm T}{\boldsymbol{\omega}}^{\star}_{\cal I}$. Solving this linear system leads to the remaining vector of the dual ${\boldsymbol{\omega}}^{\star}_{{\cal I}^{c}}\in{\mathbb{R}}^{p-m}$.
@@ Appendix E Proof of Theorem 6
The residuals of the least $\rho_{\tau}$ fit have a distribution which scale is proportional to $\xi$; so since $S^{\tau}=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}\|_{\infty}$ is a function of the dual vector which is the sign of the residuals, it is pivotal with respect to $\xi$. Moreover assuming the quantile regression estimate ${\hat{\boldsymbol{\beta}}^{\tau}_{H_{0}}}$ is asymptotically Gaussian with mean ${\boldsymbol{\beta}}^{\tau}_{H_{0}}$, we have that $X{\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}}-{\boldsymbol{Y}}=X(\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}-{\boldsymbol{\beta}_{H_{0}}^{\tau}})-X{\boldsymbol{\epsilon}}$ is asymptically pivotal with respect to ${\boldsymbol{\beta}}_{H_{0}}^{\tau}$.

### v3

@@ 5.5 Robustness of the level to non-Gaussian errors
Under the null $H_{0}:A{\boldsymbol{\beta}}={\boldsymbol{b}}$ and for a nominal level $\alpha=0.01$, we show the effective level of the $F$-test and the $\infty$-S test as a function of the degrees of freedom for Student errors d.f. $\in\{1,2,3,4,5,10,100\}$. We choose $n=100$, $p=20$, AA is the first $m=5$ rows of the finite difference matrix of Section 5.4 and $\boldsymbol{\beta}$} are $p$ samples from the standard Gaussian distribution; $\boldsymbol{b}$ is calculated as $A{\boldsymbol{\beta}}$. In a robustness analysis of the F-test, ALI96 observe that major determinant of the sensitivity to nonnormality of the errors is the extent of the nonnormality of the regressors or the extent of presence of ‘leveraged’ (influential) observations. So we simulate the entries of the matrix $X$ as a sample from $T$, where $T$ follows a student distribution with 2 degrees of freedom.
@@ 6 Conclusions
The $\infty$-Sτ test allows to test general linear null hypotheses for linear models and for any quantile $\tau$. The test function is simple to implement and can handle high-dimensional problems. We emphasize the $\infty$-S test which corresponds to the $\ell_{1}$-LASSO penalty, and showed that other norms than the $\ell_{1}$-norm could be employed, leading to other tests. SARDY2022107507 show that using $\ell_{1}$ leads to high power under sparse alternatives (that is, most entries, but not all, of $A\boldsymbol{\beta}-\boldsymbol{b}$ are zero when $H_{1}$ is true), and, under dense alternatives (that is, most or all entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ are non-zero), that the sign test based on $\ell_{2}$ has more power than the $\infty$-S test. To get nearly the best power with a single test regardless whether the alternative hypothesis is sparse or dense, SARDY2022107507 propose the $\oplus$-test.
The inversion of $\infty$-S test can be used to derive confidence regions. The $(1-\alpha)$-confidence region for $A\boldsymbol{\beta}$ based on the $\infty$-S test consists of all vectors $\boldsymbol{b}$ that are not rejected for $H_{0}:A\boldsymbol{\beta}=\boldsymbol{b}$. In particular, to test $H_{0}:\ \beta_{j}=b_{j}$ for some fixed $j\in\{1,\ldots,p\}$, one uses $A=\boldsymbol{e}_{j}^{\rm T}$, the $j$th coordinate (canonical) vector; the statistic is $s=|\boldsymbol{x}_{j}^{\rm T}\boldsymbol{\omega}|$, where $\boldsymbol{\omega}={\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j})$. So the $(1-\alpha)$-confidence interval for $\beta_{j}$ is the set of all $b_{j}$ such that $|\boldsymbol{x}_{j}^{\rm T}{\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j}(b_{j}))|\leq c_{\alpha}$, where $\hat{\boldsymbol{\beta}}_{-j}(b_{j})\in\arg\min_{\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p-1}}\|\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\tilde{{\boldsymbol{\beta}}}\|_{1}$ and $X_{-j}$ is the $X$ data matrix without its $j$th column.
@@
[…]
Acknowledgements
The research of Sylvain Sardy was partially supported by the Swiss National Science Foundation grant 200021E_213166, a partner in the Lead Agency project GAČR No. 23–06461K of the Czech Science Foundation, which supported the research of Ivan Mizera; the research of Xiaoyu Ma was supported by the National Natural Science Foundation of China(No. 12401397).
@@ Appendix A Proof of Theorem 1
Consider the primal cost function in (2.5). This primal cost being convex, the gap between the primal and dual cost functions is zero at optimality. The dual cost can be derived as follows
@@
where $\gamma()$ applied componentwise to the penalized vector ${\boldsymbol{\kappa}}=A\boldsymbol{\beta}^{\star}-\boldsymbol{b}$ is defined by $\gamma(\kappa_{k})={\rm sign}(\kappa_{k})$ if $\kappa_{k}\neq 0$, and $\gamma(0)\in[-1,1]$, for $k=1,\ldots,m$. We are interested in finding the smallest $\lambda$ for the solution to (2.5) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and (14) is
@@
[…]
@@ Appendix C Proof of Theorem 4
It is well known that the LAD optimization problem (here with $\boldsymbol{y}=\boldsymbol{y}_{-A}$ and $X=X_{-A}$) can be rewritten as the linear program: min_r_+≥0, r_-≥0,β (r_+^T, r_-^T)^T 1_2n s.t. y_-A=X_-Aβ_-A+r_+-r_-, where $\boldsymbol{1}_{2n}$ is the vector of ones of length $2n$. Using the Lagrange multiplier dual variable $\boldsymbol{\omega}$, the dual can be derived as follows.
@@
with $w_{i}=1=$ when $(r_{+})_{i}>0$, $w_{i}=-1$ when $(r_{-})_{i}>0$, so $\boldsymbol{\omega}$ is the sign of the residuals $\boldsymbol{y}_{-A}-X_{-A}\hat{{\boldsymbol{\beta}}}_{-A}=\boldsymbol{y}-X\hat{\boldsymbol{\beta}}_{H_{0}}$ when the residuals are non-zero. So the dual problems of the constrained LAD and affine LAD-LASSO when $\lambda=\lambda_{0}(\boldsymbol{Y},X,A)$ are the same.
@@ Appendix D Proof of Theorem 5
The proof follows the same lines as the proof of Theorem 1. The dual cost to (4.12) is the same except that the constraint on ${\boldsymbol{\omega}}\in[-\tau,1-\tau]^{n}$ with the same KKT conditions. We are interested in finding the smallest $\lambda$ for the solution to (4.12) to satisfy $H_{0}$. In that case, ${\boldsymbol{\kappa}}={\boldsymbol{0}}$ and ${\boldsymbol{\beta}^{\star}=\hat{\boldsymbol{\beta}}}_{H_{0}}^{\tau}$, so $\gamma(A\boldsymbol{\beta}^{\star}-\boldsymbol{b})={\cal B}^{m}_{\infty}(1)$, the unit ball of the infinite norm in ${\mathbb{R}}^{m}$, and the KKT conditions are
@@
[…]
@@ Appendix E Proof of Theorem 6
The residuals of the least $\rho_{\tau}$ fit have a distribution which scale is proportional to $\xi$; so since $S^{\tau}=\|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}\|_{\infty}$ is a function of the dual vector which is the sign of the residuals, it is pivotal with respect to $\xi$. Moreover assuming the quantile regression estimate $\hat{\boldsymbol{\beta}}^{\tau}_{H_{0}}$} is asymptotically Gaussian with mean ${\boldsymbol{\beta}}^{\tau}_{H_{0}}$, we have that $X{\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}}-{\boldsymbol{Y}}=X(\hat{\boldsymbol{\beta}}_{H_{0}}^{\tau}-{\boldsymbol{\beta}_{H_{0}}^{\tau}})-X{\boldsymbol{\epsilon}}$ is asymptotically pivotal with respect to ${\boldsymbol{\beta}}_{H_{0}}^{\tau}$.
Sylvian Sardy, Section of Mathematics, University of Geneva, rue du Conseil-Général 7-9, 1205 Geneva, Switzerland.
Ivan Mizera, Department of Probability and Mathematical Statistics, Faculty of Mathematics and Physics, Charles University Prague, Sokolovská 83, 186 75 Praha, Czechia.
Xiaoyu Ma, College of Science, National University of Defense and Technology, 1 Fuyuan Road, 410000 Changsha, China.
Hugo Gaible, Department of Mathematics, École Normale Supérieur Paris-Saclay, 4 avenue des Sciences, 91190 Gif-sur-Yvette, France.